### v3

@@ 3.1 Distance to disjoint union
    Set $\mathrm{w}(\{r,v_{j}\})=3(t-1)B$ for $j\in[n]$, and $\mathrm{w}(e)=1$ for all other edges.
@@
For $j\in[t]$, let $A_{j}=\{i\mid\exists u\in V_{i},\,\{u,v_{j}\}\in E(T)\}$. ˜1 implies that $(A_{1},\dots,A_{t})$ is a partition of $[n]$. In particular, the set of vertices of the component of $T-\{r,v_{j}\}$ containing $v_{j}$ is $\{v_{j}\}\cup\bigcup_{i\in A_{j}}V_{i}$. This implies that
@@ 3.2 Modular-width
    Add a set of vertices $Q=\{v_{j}\mid j\in[n]\}$ and all possible edges between $Q$ and $\bigcup_{i\in[3n]}V_{i}$.
@@
    Add a set of vertices $C=\{c_{z}\mid z\in[M]\}$ that form a clique and all possible edges between $C$ and $Q$, where M=4(n1)BM=4(n-1)B.
@@
This concludes the construction of $G$; we refer to Figure˜2 for an overview of the construction. We set k=M+2(n1)B=3M/2k=M+2(n-1)B=3M/2 and we claim that $\mathcal{I}$ is a yes-instance of $3$-Partition if and only if $\mathrm{stc}(G)\leq k$. Before proving this, we provide upper bounds on $G$’s modular-width and linear clique-width. Furthermore, we prove that the class of graphs produced by our reduction has shrub-depth at most $2$.
@@
  • introduce the vertices of $Q$ using label $3$,
  • join 131-3 and 232-3, and then relabel $3\to 1$,
  • introduce vertex $w$ with label $3$, and join 232-3.∎
@@
Finally, let $S=\{(1,1,1),\,(1,2,2),\,(2,1,2),\,(3,3,1),\,(2,3,1),\,(3,2,1),\,(3,4,1),\,(4,3,1)\}$. Since every graph belonging to $\mathcal{G}$ admits a tree-model of four colors and depth $2$, the statement follows. ∎
@@
$=|E(\{v_{j}\},C)|+|E(\{v_{j}\},\;\textstyle\bigcup_{i\in[3n]\setminus A_{j}}V_{i})|+|E(\textstyle\bigcup_{i\in A_{j}}V_{i},\;Q\setminus\{v_{j}\}\})|$
$=M+(n-1)B+(n-1)B=k.\qed$
@@
Next we show that for each $i\in[3n]$, there exists exactly one index $j\in[n]$ such that at least one vertex in $V_{i}$ is adjacent to $v_{j}$ in $T$. The proof is analogous to the one of ˜1.
@@
For $j\in[n]$, let $A_{j}=\{i\in[3n]\mid\exists u\in V_{i},\,\{u,v_{j}\}\in E(T)\}$. Due to ˜3 it follows that $(A_{1},\dots,A_{n})$ is a partition of $[3n]$. In particular, the set of vertices of the component of $T-\{c_{\ell},v_{j}\}$ containing $v_{j}$ is $\{v_{j}\}\cup\bigcup_{i\in A_{j}}V_{i}$. This implies that
$\mathrm{cng}_{G,T}(\{c_{\ell},v_{j}\})$ $=|E(\{v_{j}\},C)|+|E(\{v_{j}\},\;\textstyle\bigcup_{i\in[3n]\setminus A_{j}}V_{i})|+|E(\textstyle\bigcup_{i\in A_{j}}V_{i},Q\setminus\{v_{j}\})|$
$=M+\sum_{i\in[3n]\setminus V_{j}}a_{i}+(n-1)\sum_{i\in V_{j}}a_{i}$
$=M+nB+(n-2)\sum_{i\in V_{j}}a_{i}.$
Combining this with the assumption $\mathrm{cng}_{G,T}(c_{\ell},v_{j})\leq k=M+2(n-1)B$, we obtain that $\sum_{i\in V_{j}}a_{i}\leq B$. (Recall that $n\geq 3$.) Since $\sum_{i\in[3n]}a_{i}=nB$, we have $\sum_{i\in V_{j}}a_{i}=B$ for all $j\in[n]$. ∎
@@ 3.3 Maximum degree
By ˜4, at most one of $\{x_{i},z_{i}\}$ and $\{\bar{x}_{i},z_{i}\}$ may belong to $T$. Suppose to the contrary that none of $\{x_{i},z_{i}\}$ and $\{\bar{x}_{i},z_{i}\}$ belong to $T$. Let $\{y_{i},\ell_{i}\}$ and $\{\ell_{i},c_{j}\}$ be the first and second edges, respectively, in the $y_{i}$$z_{i}$ path in $T$, where $\ell_{i}\in\{x_{i},\bar{x}_{i}\}$. Observe that the edge set $\{\{x_{i},z_{i}\},\{\bar{x}_{i},z_{i}\},\{x_{i},c_{j}\}\}$ is not an edge-cut in $G$: the graph $G-\{x_{i},y_{i},\bar{x}_{i}\}$ is connected; and there is an edge not included in $\{\{x_{i},z_{i}\},\{\bar{x}_{i},z_{i}\},\{x_{i},c_{j}\}\}$ that connects $G-\{x_{i},y_{i},\bar{x}_{i}\}$ and $G[\{x_{i},y_{i},\bar{x}_{i}\}]$ (e.g., there is an edge between $x_{i}$ and $c_{j^{\prime}}$ for some $j^{\prime}\neq j$). Therefore, the edge set $E(V(T_{\{x_{i},c_{j}\},1}),V(T_{\{x_{i},c_{j}\},2}))$ is a proper super set of $\{\{x_{i},z_{i}\},\{\bar{x}_{i},z_{i}\},\{x_{i},c_{j}\}\}$. In that case,
@@
By symmetry, it suffices to show that if $\{x_{i},c_{j}\}\in E(T)$ for some $j\in[m]$, then $\{x_{i},z_{i}\}\in E(T)$. Suppose to the contrary that $\{x_{i},c_{j}\}\in E(T)$ for some $j\in[m]$ but $\{x_{i},z_{i}\}\notin E(T)$. By ˜4, $\{x_{i},y_{i}\},\{\bar{x}_{i},y_{i}\}\in E(T)$. By ˜6, $\{\bar{x}_{i},z_{i}\}\in E(T)$. Observe that the edge $\{x_{i},z_{i}\}$ and the edges incident to $c_{j}$ except for $\{x_{i},c_{j}\}$ use $\{x_{i},y_{i}\}$ in their detours in $T$. Let $j^{\prime}\in[m]$ be the index such that $j^{\prime}\neq j$ and $\{x_{i},c_{j^{\prime}}\}\in E(G)$. If $\{x_{i},c_{j^{\prime}}\}\notin E(T)$, then $\{x_{i},c_{j^{\prime}}\}$ uses $\{x_{i},y_{i}\}$ in its detour in $T$. Otherwise, the edges incident to $c_{j^{\prime}}$ except for $\{x_{i},c_{j^{\prime}}\}$ use $\{x_{i},y_{i}\}$ in their detours in $T$. In total, at least four edges not including $\{x_{i},y_{i}\}$ itself use $\{x_{i},y_{i}\}$ in their detours in $T$. Therefore, we have $\mathrm{cng}_{G,T}(\{x_{i},y_{i}\})\geq 4+\mathrm{w}_{2}(\{x_{i},y_{i}\})=4+(k-3)>k$, a contradiction.
Now we construct an assignment $\alpha\colon X\to\{0,1\}$ from $T$ by setting $\alpha(x_{i})=1$ if $\{x_{i},z_{i}\}\in E(T)$ and $\alpha(x_{i})=0$ if $\{\bar{x}_{i},z_{i}\}\in E(T)$. Note that, by ˜6, exactly one of these cases happens for each $i\in[n]$. We show that $\alpha$ is a satisfying assignment of $(X,C)$. Let $j\in[m]$. There is an index $i\in[n]$ such that $x_{i}$ or $\bar{x}_{i}$ is adjacent to $c_{j}$ in $T$. By symmetry, assume that $\{x_{i},c_{j}\}\in E(T)$. By ˜8, $\{x_{i},z_{i}\}\in E(T)$ holds. Therefore, we have $\alpha(x_{i})=1$, and thus the clause $c_{j}$ is satisfied by $\alpha$.
@@ 5.1 Distance to Clique
Let $e=\{u_{1},u_{2}\}\in E(T_{\operatorname{OPT}})$. By ˜10 and the fact that $c>2k^{3}+4k$ it follows that $\mathrm{cng}_{G,T_{\operatorname{OPT}}}(e)<2c-\frac{c}{k}+2k^{2}<2c-2k^{2}-4+2k^{2}=2(c-2)$. Assume that $|V_{u_{1}}\cap C|=x$ and $|V_{u_{2}}\cap C|=c-x$, where $x\in[2,c-2]$. In that case, it holds that $\mathrm{cng}_{G,T_{\operatorname{OPT}}}(e)\geq x\cdot(c-x)$ due to the vertices of $C$. Since $x\cdot(c-x)\geq 2\cdot(c-2)$ for all $x\in[2,c-2]$, it follows that $\mathrm{cng}_{G,T_{\operatorname{OPT}}}(e)\geq 2(c-2)$, a contradiction.
Consequently, due to ˜11 it follows that there exists a vertex $r\in V$ such that in $T_{\operatorname{OPT}}-r$, every connected component contains at most one vertex of $C$. Notice that the vertices of $S$ are present in at most $k$ such connected components. Therefore, in order to compute $T_{\operatorname{OPT}}$, we initially guess vertex $r$, and due to the indistinguishability of twin vertices, this amounts to a total of $2^{k}+k$ choices. Next, we guess the set $Q\subseteq C$, which contains the vertices of $C$ present in connected components of $T_{\operatorname{OPT}}-r$ that contain vertices of $S$. Since $|Q|\leq k$ and vertices belonging to the same twin classes are indistinguishable, there are a total of $\sum_{i=0}^{k}{(2^{k})}^{i}\leq k\cdot 2^{k^{2}}$ choices regarding the set $Q$. For the rest of the vertices $c\in C\setminus Q$, we add the edge $\{r,c\}$ in $E(T_{\operatorname{OPT}})$ (if such an edge does not exist, our guess is wrong and we discard it). Finally, since $T_{\operatorname{OPT}}$ is a tree, it suffices to guess the parent for each vertex in $S\cup Q$, among the set $S\cup Q\cup\{r\}$, which gives a total of at most ${(2k+1)}^{2k}$ possibilities. Given one such spanning tree, one can in polynomial time compute its congestion. The final running time is $2^{{\cal O}(k^{2})}n^{{\cal O}(1)}$.
@@ 5.2 Vertex integrity
    • By ˜12, it suffices to find an optimal signature of $T$ consistent with the guesses.

### v5

@@ 3.1 Distance to disjoint union
    Set $\mathrm{w}(\{r,v_{j}\})=3(t-1)B$ for $j\in[t]$, and $\mathrm{w}(e)=1$ for all other edges.
@@
For $j\in[t]$, let $A_{j}=\{i\mid\exists u\in V_{i},\,\{u,v_{j}\}\in E(T)\}$. Claim˜1 implies that $(A_{1},\dots,A_{t})$ is a partition of $[n]$. In particular, the set of vertices of the component of $T-\{r,v_{j}\}$ containing $v_{j}$ is $\{v_{j}\}\cup\bigcup_{i\in A_{j}}V_{i}$. This implies that
@@ 3.2 Modular-width
    Add a set of vertices $Q=\{v_{j}\mid j\in[n]\}$ that form a clique and all possible edges between $Q$ and $\bigcup_{i\in[3n]}V_{i}$.
@@
    Add a set of vertices $C=\{c_{z}\mid z\in[M]\}$ that form a clique and all possible edges between $C$ and $Q$, where M=4(n1)B+2(n1)M=4(n-1)B+2(n-1).
@@
This concludes the construction of $G$; we refer to Figure˜2 for an overview of the construction. We set k=M+2(n1)B+(n1)=3M/2k=M+2(n-1)B+(n-1)=3M/2 and we claim that $\mathcal{I}$ is a yes-instance of $3$-Partition if and only if $\mathrm{stc}(G)\leq k$. Before proving this, we provide upper bounds on $G$’s modular-width and linear clique-width. Furthermore, we prove that the class of graphs produced by our reduction has shrub-depth at most $2$.
@@
  • introduce vertex $w$ with label $3$, and join 2233,
  • relabel $2\to 1$,
  • for each vertex $v\in Q$, (i) introduce $v$ using label $2$; (ii) join 1122; (iii) relabel $2\to 1$.∎
@@
Finally, let $S=\{(1,1,1),\,(1,2,2),\,(2,1,2),\,(2,2,1),\,(3,3,1),\,(2,3,1),\,(3,2,1),\,(3,4,1),\,(4,3,1)\}$. Since every graph belonging to $\mathcal{G}$ admits a tree-model of four colors and depth $2$, the statement follows. ∎
@@
$=|E(\{v_{j}\},C)|+|E(\{v_{j}\},\;\textstyle\bigcup_{i\in[3n]\setminus A_{j}}V_{i})|+|E(\textstyle\{v_{j}\}\cup\bigcup_{i\in A_{j}}V_{i},\;Q\setminus\{v_{j}\}\})|$
$=M+(n-1)B+(n-1)+(n-1)B=k.\qed$
@@
Next we show that for each $i\in[3n]$, there exists exactly one index $j\in[n]$ such that at least one vertex in $V_{i}$ is adjacent to $v_{j}$ in $T$. The proof is analogous to the one of Claim˜1.
@@
For $j\in[n]$, let $A_{j}=\{i\in[3n]\mid\exists u\in V_{i},\,\{u,v_{j}\}\in E(T)\}$. Due to Claim˜3 it follows that $(A_{1},\dots,A_{n})$ is a partition of $[3n]$. In particular, the set of vertices of the component of $T-\{c_{\ell},v_{j}\}$ containing $v_{j}$ is $\{v_{j}\}\cup\bigcup_{i\in A_{j}}V_{i}$. This implies that
$\mathrm{cng}_{G,T}(\{c_{\ell},v_{j}\})$ $=|E(\{v_{j}\},C)|+|E(\{v_{j}\},\;\textstyle\bigcup_{i\in[3n]\setminus A_{j}}V_{i})|+|E(\textstyle\{v_{j}\}\cup\bigcup_{i\in A_{j}}V_{i},Q\setminus\{v_{j}\})|$
$=M+\sum_{i\in[3n]\setminus V_{j}}a_{i}+(n-1)+(n-1)\sum_{i\in V_{j}}a_{i}$
$=M+(n-1)+nB+(n-2)\sum_{i\in V_{j}}a_{i}.$
Combining this with the assumption $\mathrm{cng}_{G,T}(c_{\ell},v_{j})\leq k=M+(n-1)+2(n-1)B$, we obtain that $\sum_{i\in V_{j}}a_{i}\leq B$. (Recall that $n\geq 3$.) Since $\sum_{i\in[3n]}a_{i}=nB$, we have $\sum_{i\in V_{j}}a_{i}=B$ for all $j\in[n]$. ∎
@@ 3.3 Maximum degree
By Claim˜4, at most one of $\{x_{i},z_{i}\}$ and $\{\bar{x}_{i},z_{i}\}$ may belong to $T$. Suppose to the contrary that none of $\{x_{i},z_{i}\}$ and $\{\bar{x}_{i},z_{i}\}$ belong to $T$. Let $\{y_{i},\ell_{i}\}$ and $\{\ell_{i},c_{j}\}$ be the first and second edges, respectively, in the $y_{i}$$z_{i}$ path in $T$, where $\ell_{i}\in\{x_{i},\bar{x}_{i}\}$. Observe that the edge set $\{\{x_{i},z_{i}\},\{\bar{x}_{i},z_{i}\},\{x_{i},c_{j}\}\}$ is not an edge-cut in $G$: the graph $G-\{x_{i},y_{i},\bar{x}_{i}\}$ is connected; and there is an edge not included in $\{\{x_{i},z_{i}\},\{\bar{x}_{i},z_{i}\},\{x_{i},c_{j}\}\}$ that connects $G-\{x_{i},y_{i},\bar{x}_{i}\}$ and $G[\{x_{i},y_{i},\bar{x}_{i}\}]$ (e.g., there is an edge between $x_{i}$ and $c_{j^{\prime}}$ for some $j^{\prime}\neq j$). Therefore, the edge set $E(V(T_{\{x_{i},c_{j}\},1}),V(T_{\{x_{i},c_{j}\},2}))$ is a proper super set of $\{\{x_{i},z_{i}\},\{\bar{x}_{i},z_{i}\},\{x_{i},c_{j}\}\}$. In that case,
@@
By symmetry, it suffices to show that if $\{x_{i},c_{j}\}\in E(T)$ for some $j\in[m]$, then $\{x_{i},z_{i}\}\in E(T)$. Suppose to the contrary that $\{x_{i},c_{j}\}\in E(T)$ for some $j\in[m]$ but $\{x_{i},z_{i}\}\notin E(T)$. By Claim˜4, $\{x_{i},y_{i}\},\{\bar{x}_{i},y_{i}\}\in E(T)$. By Claim˜6, $\{\bar{x}_{i},z_{i}\}\in E(T)$. Observe that the edge $\{x_{i},z_{i}\}$ and the edges incident to $c_{j}$ except for $\{x_{i},c_{j}\}$ use $\{x_{i},y_{i}\}$ in their detours in $T$. Let $j^{\prime}\in[m]$ be the index such that $j^{\prime}\neq j$ and $\{x_{i},c_{j^{\prime}}\}\in E(G)$. If $\{x_{i},c_{j^{\prime}}\}\notin E(T)$, then $\{x_{i},c_{j^{\prime}}\}$ uses $\{x_{i},y_{i}\}$ in its detour in $T$. Otherwise, the edges incident to $c_{j^{\prime}}$ except for $\{x_{i},c_{j^{\prime}}\}$ use $\{x_{i},y_{i}\}$ in their detours in $T$. In total, at least four edges not including $\{x_{i},y_{i}\}$ itself use $\{x_{i},y_{i}\}$ in their detours in $T$. Therefore, we have $\mathrm{cng}_{G,T}(\{x_{i},y_{i}\})\geq 4+\mathrm{w}_{2}(\{x_{i},y_{i}\})=4+(k-3)>k$, a contradiction.
Now we construct an assignment $\alpha\colon X\to\{0,1\}$ from $T$ by setting $\alpha(x_{i})=1$ if $\{x_{i},z_{i}\}\in E(T)$ and $\alpha(x_{i})=0$ if $\{\bar{x}_{i},z_{i}\}\in E(T)$. Note that, by Claim˜6, exactly one of these cases happens for each $i\in[n]$. We show that $\alpha$ is a satisfying assignment of $(X,C)$. Let $j\in[m]$. There is an index $i\in[n]$ such that $x_{i}$ or $\bar{x}_{i}$ is adjacent to $c_{j}$ in $T$. By symmetry, assume that $\{x_{i},c_{j}\}\in E(T)$. By Claim˜8, $\{x_{i},z_{i}\}\in E(T)$ holds. Therefore, we have $\alpha(x_{i})=1$, and thus the clause $c_{j}$ is satisfied by $\alpha$.
@@ 5.1 Distance to Clique
Let $e=\{u_{1},u_{2}\}\in E(T_{\operatorname{OPT}})$. By Claim˜10 and the fact that $c>2k^{3}+4k$ it follows that $\mathrm{cng}_{G,T_{\operatorname{OPT}}}(e)<2c-\frac{c}{k}+2k^{2}<2c-2k^{2}-4+2k^{2}=2(c-2)$. Assume that $|V_{u_{1}}\cap C|=x$ and $|V_{u_{2}}\cap C|=c-x$, where $x\in[2,c-2]$. In that case, it holds that $\mathrm{cng}_{G,T_{\operatorname{OPT}}}(e)\geq x\cdot(c-x)$ due to the vertices of $C$. Since $x\cdot(c-x)\geq 2\cdot(c-2)$ for all $x\in[2,c-2]$, it follows that $\mathrm{cng}_{G,T_{\operatorname{OPT}}}(e)\geq 2(c-2)$, a contradiction.
Consequently, due to Claim˜11 it follows that there exists a vertex $r\in V$ such that in $T_{\operatorname{OPT}}-r$, every connected component contains at most one vertex of $C$. Notice that the vertices of $S$ are present in at most $k$ such connected components. Therefore, in order to compute $T_{\operatorname{OPT}}$, we initially guess vertex $r$, and due to the indistinguishability of twin vertices, this amounts to a total of $2^{k}+k$ choices. Next, we guess the set $Q\subseteq C$, which contains the vertices of $C$ present in connected components of $T_{\operatorname{OPT}}-r$ that contain vertices of $S$. Since $|Q|\leq k$ and vertices belonging to the same twin classes are indistinguishable, there are a total of $\sum_{i=0}^{k}{(2^{k})}^{i}\leq k\cdot 2^{k^{2}}$ choices regarding the set $Q$. For the rest of the vertices $c\in C\setminus Q$, we add the edge $\{r,c\}$ in $E(T_{\operatorname{OPT}})$ (if such an edge does not exist, our guess is wrong and we discard it). Finally, since $T_{\operatorname{OPT}}$ is a tree, it suffices to guess the parent for each vertex in $S\cup Q$, among the set $S\cup Q\cup\{r\}$, which gives a total of at most ${(2k+1)}^{2k}$ possibilities. Given one such spanning tree, one can in polynomial time compute its congestion. The final running time is $2^{{\cal O}(k^{2})}n^{{\cal O}(1)}$.
@@ 5.2 Vertex integrity
    • By Claim˜12, it suffices to find an optimal signature of $T$ consistent with the guesses.